A painting canvas has a length of 6 inches longer than its width. If the area is 91 inches, what are the dimensions of the canvas?

Answer by checkley77(12844)

L=W+6
A=LW
91=(W+6)W
91=W^2+6W
W^2+6W-91=0
(W+13)(W-7)=0
W-7=0
W=7 ANS FOR THE WIDTH.
L=7+6=13 ANS. FOR THE LENGTH.
PROOF:
91=13*7
91=91